# Multimodal Medical Code Tokenizer

### v2

@@ Abstract
Foundation models trained on patient electronic health records (EHRs) require tokenizing medical data into sequences of discrete vocabulary items. Existing tokenizers treat medical codes from EHRs as isolated textual tokens. However, each medical code is defined by its textual description, its position in ontological hierarchies, and its relationships to other codes, such as disease co-occurrences and drug-treatment associations. Medical vocabularies contain more than 600,000 codes with critical information for clinical reasoning. We introduce MedTok, a multimodal medical code tokenizer that uses the text descriptions and relational context of codes. MedTok processes text using a language model encoder and encodes the relational structure with a graph encoder. It then quantizes both modalities into a unified token space, preserving modality-specific and cross-modality information. We integrate MedTok into five EHR models and evaluate it on operational and clinical tasks across in-patient and out-patient datasets, including outcome prediction, diagnosis classification, drug recommendation, and risk stratification. Swapping standard EHR tokenizers with MedTok improves AUPRC across all EHR models, by 4.10% on MIMIC-III, 4.78% on MIMIC-IV, and 11.30% on EHRShot, with the largest gains in drug recommendation. Beyond EHR modeling, we demonstrate using MedTok tokenizer with medical QA systems. Our results demonstrate the potential of MedTok as a unified tokenizer for medical codes, improving tokenization for medical foundation models.
Medical Code, Tokenizer
@@ 1 Introduction
Figure 1: MedTok is a multimodal tokenizer for medical codes that combines text embeddings from code descriptions with graph-based representations of dependencies from ontologies and terminologies. It is a general tokenizer that sheds light on how optimized tokenization benefits transformer models in medicine.
Electronic health records (EHRs) are the backbone of modern healthcare, capturing a person’s health state with increasing precision across diverse modalities. Structured EHR data, encoded through standardized medical codes, support a wide range of applications, from personalized risk prediction (Goldstein et al., 2016; Yu et al., 2024b) and disease trajectory modeling (Jensen et al., 2017; Heumos et al., 2024) to emulation of clinical trials (Katsoulakis et al., 2024; Kraljevic et al., 2024). The cornerstone of structured EHRs is medical coding systems, which assign standardized alphanumeric codes to various aspects of patient health, including diseases, procedures, medications, and laboratory tests. These codes come from widely used terminologies such as ICD-9, ICD-10, SNOMED CT, CPT, and ATC, among others (Foley et al., 1992; Organization et al., 1988; Organization, 2004; Donnelly et al., 2006; Dotson, 2013; Miller & Britt, 1995). Although essential for interoperability, medical codes introduce challenges for models, particularly in the tokenization process, which transforms structured EHR data into token sequences that foundation models can process.
@@
Several models attempt to enrich the representations of medical codes by incorporating external knowledge from LLM (Jiang et al., 2023b; Zhu et al., 2024; Xu et al., 2024). Methods like GraphCare and RAW-EHR prompt LLMs to generate structured knowledge triplets of medical codes. Although effective in specific tasks, these approaches suffer from limited generalizability, a heavy reliance on knowledge generated by LLM, and a lack of a unified framework for handling various medical coding systems. Despite advances in medical representation learning, a unified tokenizer that integrates textual and structured relational knowledge across coding systems remains an open challenge.
Present work. We introduce MedTok, a multimodal medical code tokenizer that integrates textual descriptions and graph-based dependencies from biomedical ontologies (Figure 1). Unlike standard tokenization methods that treat medical codes as isolated textual tokens, MedTok captures both semantic meaning and structured relationships by encoding multiple modalities into a unified token space. MedTok operates in three stages. Multimodal encoding first extracts text embeddings from medical code descriptions and graph-based representations from biomedical knowledge graphs using separate encoders. Next, vector quantization maps both modalities into a shared token space, generating distinct text-informed and graph-informed token embeddings while preserving cross-modality interactions. Finally, optimization for expressivity ensures that token representations capture hierarchical relationships, semantic equivalence across different coding systems, and dependencies such as comorbidities and drug interactions.
We integrate MedTok into five EHR models and evaluate it in clinical and operational tasks that span the inpatient (MIMIC-III, MIMIC-IV) and outpatient (EHRShot) settings. These tasks include disease prediction, operational outcome modeling, drug recommendation, patient risk stratification, and operational outcomes. Key contributions:
@@ 2 Related work
Vector-Quantized Tokenizers. Tokenization strategies often vary according to the problem domain and data modality where recent work has highlighted the benefits of discrete tokenization (Du et al., 2024). This process involves partitioning the input according to a finite set of tokens, often held in a codebook (this concept is independent of medical coding despite the similar name), and the quantization process involves learning a mapping from input data to the optimal set of tokens according to a pre-defined objective such as reconstruction loss (Van Den Oord et al., 2017).
Recent work has highlighted the ability of vector quantized (VQ-based) tokenization to effectively compress semantic information(Gu et al., 2024). This approach is particularly successful for tokenizing inputs with an inherent semantic structure such as graphs (Yang et al., 2023a; Wang et al., 2024c), speech (Zeghidour et al., 2021; Baevski et al., 2019), and time (Yu et al., 2021) as well as complex tasks like recommendation retrieval (Wang et al., 2024d; Rajput et al., 2023; Sun et al., 2024) and image synthesis (Zhang et al., 2023; Yu et al., 2021).
Another significant advantage to VQ-based tokenization is the natural integration of multiple modalities. By learning a shared latent space across modalities, each modality can jointly modeled using a common token vocabulary (Agarwal et al., 2025; Yu et al., 2023). TokenFlow leverages a dual-codebook design that allows for correlations across modalities through a dual encoder (Qu et al., 2024).
@@
Structured EHR, transformer-based, and foundation models. Structured EHR models leverage patient records to learn representations for clinical prediction and operational healthcare tasks. These models differ from medical LLMs (Singhal et al., 2025; Tu et al., 2024; Singhal et al., 2023), which are typically trained on free-text clinical notes (Jiang et al., 2023a) and biomedical literature rather than structured EHR data. BEHRT (Li et al., 2020) applies deep bidirectional learning to predict future medical events, encoding disease codes, age, and visit sequences using self-attention. TransformEHR (Yang et al., 2023b) adopts an encoder-decoder transformer with visit-level masking to pretrain on EHRs, enabling multi-task prediction. GT-BEHRT (Poulain & Beheshti, 2024) models intra-visit dependencies as a graph, using a graph transformer to learn visit representations before processing patient-level sequences with a transformer encoder. Other models enhance EHR representations with external knowledge. GraphCare (Jiang et al., 2023b) integrates large language models and biomedical knowledge graphs to construct patient-specific graphs processed via a Bi-attention Augmented Graph Neural Network. Mult-EHR (Chan et al., 2024) introduces multi-task heterogeneous graph learning with causal denoising to address data heterogeneity and confounding effects. ETHOS (Renc et al., 2024) tokenizes patient health timelines for transformer-based pretraining, achieving zero-shot performance. While these models focus on learning patient representations, MedTok serves a different role as a medical code tokenizer. It can be integrated into any structured EHR, transformer-based, or other foundation model, improving how medical codes are tokenized before being processed. Unlike these models, which rely on predefined tokenization schemes, MedTok optimizes the tokenization process itself.
@@ 3 Approach
Problem definition. We formulate our problem as follows. Our goal is to train a multimodal tokenizer $f(\cdot)$ so that $\mathcal{T}=f(\mathcal{D}(m),\mathcal{G}(m))$, where $\mathcal{T}=[t_{1},t_{2},...,t_{T}]$ and $t_{i}\in\mathcal{V},1\leq i\leq T$. Then the generated $\mathcal{T}$ for medical code $m$ could be integrated to any EHR-based models $h(\cdot)$ and LMs models $p(\cdot)$ to perform predictive or generative tasks.
Overview. Fig. 2 illustrates the architecture of MedTok, which takes both the medical code description and contextual knowledge from biomedical KGs as input. MedTok takes two steps, multimodal tokenization and token packing.
@@ 3.1 Multimodal Tokenization
Given a medical code $m$, paired with its description $t$ and its biological subgraph $G$, MedTok first adopts the text encoder, denoted as $\mathrm{E}_{t}$ and the graph encoder, denoted as $\mathrm{E}_{g}$, to generate two embeddings: the text semantic embedding $\mathbf{x}_{t}\in\mathbb{R}^{d_{t}}$ for $t$ and the graph-level embedding $\mathbf{x}_{g}\in\mathbb{R}^{d_{g}}$ for $G$. These embeddings are computed as $\mathbf{x}_{t}=\mathrm{E}_{t}(t)$ and $\mathbf{x}_{g}=\mathrm{E}_{g}(G)$ for $G$, where $\mathrm{E}_{t}$ and $\mathrm{E}_{g}$ represent the text and graph encoders, respectively.
Modality-specific Embeddings. MedTok then adopts two linear projectors: $f_{t}:\mathbb{R}^{d_{t}}\rightarrow\mathbb{R}^{d}$ and $f_{g}:\mathbb{R}^{d_{g}}\rightarrow\mathbb{R}^{d}$, to generate modality-specific embeddings $\mathbf{e}_{t}^{s}\in\mathbb{R}^{d}$ and $\mathbf{e}_{g}^{s}\in\mathbb{R}^{d}$, respectively, where $\mathbf{e}_{t}^{s}=f_{t}(\mathbf{x}_{t})$, $\mathbf{e}_{g}^{s}=f_{g}(\mathbf{x}_{g})$, and $d$ is the dimension of specific embeddings.
Cross-modality Embeddings. Moreover, MedTok incorporates a cross-attention module to derive cross-modality embeddings $\mathbf{e}_{t}^{c}\in\mathbb{R}^{d}$ and $\mathbf{e}_{g}^{c}\in\mathbb{R}^{d}$, Specifically, the embedding $\mathbf{e}_{t}^{c}$ is computed as:
$\mathbf{e}_{t}^{c}=\mathrm{softmax}\left(\frac{\mathbf{W}_{q}\mathbf{x}_{t}(\mathbf{W}_{k}\mathbf{x}_{g})^{T}}{\sqrt{d}}\right)(\mathbf{W}_{v}\mathbf{x}_{g})$ (1)
Similarly, the embedding $\mathbf{e}_{g}^{c}$ is given by:
$\mathbf{e}_{g}^{c}=\mathrm{softmax}\left(\frac{\mathbf{W}_{q}\mathbf{x}_{g}(\mathbf{W}_{k}\mathbf{x}_{t})^{T}}{\sqrt{d}}\right)(\mathbf{W}_{v}\mathbf{x}_{t})$ (2)
where $\mathbf{W}_{q},\mathbf{W}_{k},\mathbf{W}_{v}\in\mathbb{R}^{d\times d}$ represents the query, key, and value weight matrix.
Tokenization. After generating modality-specific and cross-modality embeddings, for each embedding, MedTok quantizes the embedding into $K$ tokens by querying a unified codebook $\mathbf{C}\in\mathbb{R}^{N\times d}$. The $K$ tokens are identified by the top $K$ nearest vectors in the codebook.
In detail, for any modality-specific or cross-modality embedding $\mathbf{e}_{:}$, its quantized tokens $\mathcal{I}(\mathbf{e}_{:})$ is formulated by:
@@
where $\text{dist}(:,:)$ denotes the Euclidean distance, $|\mathcal{I}(\mathbf{e}_{:})|=K$, and $\mathbf{C}_{i}=\mathbf{C}[i,:]$. Then MedTok assigns a weight to each token $k\in\mathcal{I}(\mathbf{e}_{:})$ based on the distance between $\mathbf{e}_{:}$ and its corresponding vector $\mathbf{C}_{k}=\mathbf{C}[k,:]$. These weighted tokens are then summed together to obtain the quantized vector for $\mathbf{e}_{:}$, denoted as $\mathbf{\hat{e}}_{:}$, which is given by:
@@
Following vector quantization conventions, we employ a straight-through gradient estimator: $\mathbf{e}_{:}=\mathrm{sg}[\mathbf{e}_{:}-\mathbf{\hat{e}}_{:}]+\mathbf{\hat{e}}_{:}$ where $\mathbf{sg}[\cdot]$ denotes the stop-gradient operation. The codebook learning objective is $\mathcal{L}(\mathbf{e}_{:},\mathbf{\hat{e}}_{:})=\|\mathrm{sg}[\mathbf{\hat{e}}_{:}]-\mathbf{e}_{:}\|_{2}^{2}+\alpha*\|\mathbf{\hat{e}}_{:}-\mathrm{sg}[\mathbf{e}_{:}]\|_{2}^{2}$, where $\alpha$ is a hyperparameter.
To preserve the distinctiveness of modality-specific and cross-modality embeddings, MedTok divides the entire codebook into three regions: a text-specific region, a graph-specific region, and a shared region, and then queries distinct regions of the codebook to generate their respective tokens and quantized vectors, which are represented by: ($\mathcal{I}(\mathbf{e}_{t}^{s})$, $\mathcal{I}(\mathbf{e}_{g}^{s})$, $\mathcal{I}(\mathbf{e}_{t}^{c})$, $\mathcal{I}(\mathbf{e}_{g}^{c})$) and ($\mathbf{\hat{e}}_{t}^{s}$, $\mathbf{\hat{e}}_{g}^{s}$, $\mathbf{\hat{e}}_{t}^{c}$, $\mathbf{\hat{e}}_{g}^{c}$). Consequently, the codebook learning objective is rewritten by:
$\mathcal{L}_{C}=\mathcal{L}(\mathbf{e}_{t}^{s},\mathbf{\hat{e}}_{t}^{s})+\mathcal{L}(\mathbf{e}_{g}^{s},\mathbf{\hat{e}}_{g}^{s})+\mathcal{L}(\mathbf{e}_{t}^{c},\mathbf{\hat{e}}_{t}^{c})+\mathcal{L}(\mathbf{e}_{g}^{c},\mathbf{\hat{e}}_{g}^{c})$ (5)

### v3

@@ Abstract
Foundation models trained on patient electronic health records (EHRs) require tokenizing medical data into sequences of discrete vocabulary items. Existing tokenizers treat medical codes from EHRs as isolated textual tokens. However, each medical code is defined by its textual description, its position in ontological hierarchies, and its relationships to other codes, such as disease co-occurrences and drug-treatment associations. Medical vocabularies contain more than 600,000 codes with critical information for clinical reasoning. We introduce MedTok, a multimodal medical code tokenizer that uses the text descriptions and relational context of codes. MedTok processes text using a language model encoder and encodes the relational structure with a graph encoder. It then quantizes both modalities into a unified token space, preserving modality-specific and cross-modality information. We integrate MedTok into five EHR models and evaluate it on operational and clinical tasks across in-patient and out-patient datasets, including outcome prediction, diagnosis classification, drug recommendation, and risk stratification. Swapping standard EHR tokenizers with MedTok improves AUPRC across all EHR models, by 4.10% on MIMIC-III, 4.78% on MIMIC-IV, and 11.32% on EHRShot, with the largest gains in drug recommendation. Beyond EHR modeling, we demonstrate using MedTok tokenizer with medical QA systems. Our results demonstrate the potential of MedTok as a unified tokenizer for medical codes, improving tokenization for medical foundation models.
Medical Code, Tokenizer, EHR
Figure 1: MedTok is a multimodal tokenizer of medical codes that combines text descriptions of codes with relational representation of dependencies between codes in clinical ontologies and medical terminologies. MedTok is a general-purpose tokenizer that can be integrated into any transformer-based model or system that requires tokenization.
@@ 1 Introduction
Electronic health records (EHRs) are the backbone of modern healthcare, capturing a person’s health state with increasing precision across diverse modalities. Structured EHR data, encoded through standardized medical codes, support a wide range of applications, from personalized risk prediction (Goldstein et al., 2016; Yu et al., 2024b) and disease trajectory modeling (Jensen et al., 2017; Heumos et al., 2024) to emulation of clinical trials (Katsoulakis et al., 2024; Kraljevic et al., 2024). The cornerstone of structured EHRs is medical coding systems, which assign standardized alphanumeric codes to various aspects of patient health, including diseases, procedures, medications, and laboratory tests. These codes come from widely used terminologies such as ICD-9, ICD-10, SNOMED CT, CPT, and ATC, among others (Foley et al., 1992; Organization et al., 1988; Organization, 2004; Donnelly et al., 2006; Dotson, 2013; Miller & Britt, 1995). Medical codes are essential for interoperability across healthcare systems but complicate tokenization, the process converting structured EHR data into token sequences for model training.
@@
Several models attempt to enrich the representations of medical codes by incorporating external knowledge from Large Language Models (LLMs) (Jiang et al., 2023b; Zhu et al., 2024; Xu et al., 2024). Methods like GraphCare and RAM-EHR prompt LLMs to generate structured knowledge triplets of medical codes or summarize retrieved knowledge. Although effective in specific tasks, these approaches suffer from limited generalizability, a heavy reliance on knowledge generated by LLMs, and a lack of a unified framework for handling various medical coding systems. Despite advances in medical representation learning, a unified tokenizer that integrates textual and structured relational knowledge across coding systems remains an open challenge.
Present work. We introduce MedTok (https://github.com/mims-harvard/MedTok), a multimodal medical code tokenizer that integrates textual descriptions and graph-based dependencies from biomedical ontologies (Figure 1). Unlike standard tokenization methods that treat medical codes as isolated textual tokens, MedTok captures both semantic meaning and structured relationships by encoding multiple modalities into a unified token space. MedTok operates in three stages. Multimodal encoding first extracts text embeddings from medical code descriptions and graph-based representations from biomedical knowledge graphs using separate encoders. Next, vector quantization maps both modalities into a shared token space, generating distinct text-informed and graph-informed token embeddings while preserving cross-modality interactions. Finally, optimization for expressivity ensures that token representations capture hierarchical relationships, semantic equivalence across different coding systems, and dependencies such as comorbidities and drug interactions.
We integrate MedTok into five EHR models and evaluate it in clinical and operational tasks in inpatient (MIMIC-III, MIMIC-IV) and outpatient (EHRShot) settings. These tasks include disease prediction, operational outcome modeling, drug recommendation, patient risk stratification, and operational outcomes. Our contributions are:
@@ 2 Related work
Vector-quantized tokenizers. Tokenization strategies often vary according to the problem domain and data modality where recent work has highlighted the benefits of discrete tokenization (Du et al., 2024). This process involves partitioning the input according to a finite set of tokens, often held in a codebook (this concept is independent of medical coding despite the similar name), and the quantization process involves learning a mapping from input data to the optimal set of tokens according to a pre-defined objective such as reconstruction loss (Van Den Oord et al., 2017). Recent work has highlighted the ability of vector quantized (VQ-based) tokenization to effectively compress semantic information (Gu et al., 2024). This approach is particularly successful for tokenizing inputs with an inherent semantic structure such as graphs (Yang et al., 2023a; Wang et al., 2024c), speech (Zeghidour et al., 2021; Baevski et al., 2019), and time (Yu et al., 2021) as well as complex tasks like recommendation retrieval (Wang et al., 2024d; Rajput et al., 2023; Sun et al., 2024) and image synthesis (Zhang et al., 2023; Yu et al., 2021). Another significant advantage to VQ-based tokenization is the natural integration of multiple modalities. By learning a shared latent space across modalities, each modality can jointly modeled using a common token vocabulary (Agarwal et al., 2025; Yu et al., 2023). TokenFlow leverages a dual-codebook design that allows for correlations across modalities through a dual encoder (Qu et al., 2024).
@@
Structured EHR, transformer-based, and foundation models. Structured EHR models leverage patient records to learn representations for clinical prediction and operational healthcare tasks. These models differ from medical LLMs (Singhal et al., 2025; Tu et al., 2024; Singhal et al., 2023), which are typically trained on free-text clinical notes (Jiang et al., 2023a) and biomedical literature rather than structured EHR data. BEHRT (Li et al., 2020) applies deep bidirectional learning to predict future medical events, encoding disease codes, age, and visit sequences using self-attention. TransformEHR (Yang et al., 2023b) adopts an encoder-decoder transformer with visit-level masking to pretrain on EHRs, enabling multi-task prediction. GT-BEHRT (Poulain & Beheshti, 2024) models intra-visit dependencies as a graph, using a graph transformer to learn visit representations before processing patient-level sequences with a transformer encoder. Other models enhance EHR representations with external knowledge. GraphCare (Jiang et al., 2023b) integrates LLMs and biomedical knowledge graphs to construct patient-specific graphs processed via a Bi-attention Augmented Graph Neural Network. MulT-EHR (Chan et al., 2024) introduces multi-task heterogeneous graph learning with causal denoising to address data heterogeneity and confounding effects. ETHOS (Renc et al., 2024) tokenizes patient health timelines for transformer-based pretraining, achieving zero-shot performance. While these models focus on learning patient representations, MedTok serves a different role as a medical code tokenizer. It can be integrated into any structured EHR, transformer-based, or other foundation model, improving how medical codes are tokenized before being processed. Unlike these models, which rely on predefined tokenization schemes, MedTok optimizes the tokenization process itself.
@@ 3 Approach
Problem definition. Our goal is to train a multimodal tokenizer $f(\cdot)$ so that $\mathcal{T}=f(\mathcal{D}(m),\mathcal{G}(m))$, where $\mathcal{T}=[t_{1},t_{2},...,t_{T}]$ and $t_{i}\in\mathcal{V},1\leq i\leq T$. Then the generated $\mathcal{T}$ for medical code $m$ could be integrated to any EHR-based models $h(\cdot)$ and LMs or LLMs $p(\cdot)$ to perform predictive or generative tasks.
Figure 2 illustrates the architecture of MedTok, which takes both the medical code description and contextual knowledge from biomedical KGs as input. MedTok takes two steps, multimodal tokenization and token packing.
@@ 3.1 Multimodal tokenization
Given a medical code $m$, paired with its description $t$ and its biological subgraph $G$, MedTok first adopts the text encoder, denoted as $\mathrm{E}_{t}$ and the graph encoder, denoted as $\mathrm{E}_{g}$, to generate two embeddings: the text semantic embedding $\mathbf{x}_{t}\in\mathbb{R}^{d_{t}}$ for $t$ and the graph-level embedding $\mathbf{x}_{g}\in\mathbb{R}^{d_{g}}$ for $G$. These embeddings are computed as $\mathbf{x}_{t}=\mathrm{E}_{t}(t)$ and $\mathbf{x}_{g}=\mathrm{E}_{g}(G)$, where $\mathrm{E}_{t}$ and $\mathrm{E}_{g}$ represent the text and graph encoders, respectively.
Modality-specific information. MedTok adopts two linear projectors: $f_{t}:\mathbb{R}^{d_{t}}\rightarrow\mathbb{R}^{d}$ and $f_{g}:\mathbb{R}^{d_{g}}\rightarrow\mathbb{R}^{d}$, to generate text and graph modality-specific embeddings, $\mathbf{e}_{t}^{s}\in\mathbb{R}^{d}$ and $\mathbf{e}_{g}^{s}\in\mathbb{R}^{d}$, respectively, where $\mathbf{e}_{t}^{s}=f_{t}(\mathbf{x}_{t})$, $\mathbf{e}_{g}^{s}=f_{g}(\mathbf{x}_{g})$, and $d$ is the dimension of specific embeddings.
Cross-modality information. MedTok incorporates a cross-attention module to specify cross-modality embeddings $\mathbf{e}_{t}^{c}\in\mathbb{R}^{d}$ and $\mathbf{e}_{g}^{c}\in\mathbb{R}^{d}$. Here, embedding $\mathbf{e}_{t}^{c}$ is generated as:
$\mathbf{e}_{t}^{c}=\mathrm{softmax}\left(\frac{\mathbf{W}_{q}^{t}\mathbf{x}_{t}(\mathbf{W}_{k}^{g}\mathbf{x}_{g})^{T}}{\sqrt{d}}\right)(\mathbf{W}_{v}^{g}\mathbf{x}_{g})$ (1)
where $\mathbf{W}_{q}^{t}\in\mathbb{R}^{d\times d_{t}},\mathbf{W}_{k}^{g}\in\mathbb{R}^{d\times d_{g}}$, and $\mathbf{W}_{v}^{g}\in\mathbb{R}^{d\times d_{g}}$ represent the query, key and value weight matrices. Similarly, the embedding $\mathbf{e}_{g}^{c}$ is given by:
$\mathbf{e}_{g}^{c}=\mathrm{softmax}\left(\frac{\mathbf{W}_{q}^{g}\mathbf{x}_{g}(\mathbf{W}_{k}^{t}\mathbf{x}_{t})^{T}}{\sqrt{d}}\right)(\mathbf{W}_{v}^{t}\mathbf{x}_{t})$ (2)
where $\mathbf{W}_{q}^{g}\in\mathbb{R}^{d\times d_{g}},\mathbf{W}_{k}^{t}\in\mathbb{R}^{d\times d_{t}}$, and $\mathbf{W}_{v}^{t}\in\mathbb{R}^{d\times d_{t}}$ represents the query, key, and value weight matrices.
Tokenization. After generating modality-specific and cross-modality embeddings, MedTok quantizes them into $K$ tokens using a unified $codebook$ $\mathbf{C}\in\mathbb{R}^{N\times d}$. The tokens are selected as the top $K$ nearest vectors in $\mathbf{C}$.
Given a modality-specific or cross-modality embedding $\mathbf{e}_{:}$, its quantized tokens $\mathcal{I}(\mathbf{e}_{:})$ are given by:
@@
where $\text{dist}(:,:)$ denotes Euclidean distance, $|\mathcal{I}(\mathbf{e}_{:})|=K$, and $\mathbf{C}_{i}=\mathbf{C}[i,:]$. Then MedTok assigns a weight to each token $k\in\mathcal{I}(\mathbf{e}_{:})$ based on distance between $\mathbf{e}_{:}$ and its corresponding vector $\mathbf{C}_{k}=\mathbf{C}[k,:]$. These weighted tokens are aggregated into a quantized vector denoted as $\mathbf{\hat{e}}_{:}$, which is given by:
@@
We use a straight-through gradient estimator: $\mathbf{\hat{e}}_{:}\leftarrow\mathrm{sg}[\mathbf{\hat{e}}_{:}-\mathbf{e}_{:}]+\mathbf{e}_{:}$ where $\mathrm{sg}[\cdot]$ denotes the stop-gradient operation commonly used for vector quantization. The $codebook$ learning objective is: $\mathcal{L}(\mathbf{e}_{:},\mathbf{\hat{e}}_{:})=\|\mathrm{sg}[\mathbf{e}_{:}]-\mathbf{\hat{e}}_{:}\|_{2}^{2}+\alpha\|\mathbf{e}_{:}-\mathrm{sg}[\mathbf{\hat{e}}_{:}]\|_{2}^{2}$, where the second term represents commitment loss with a balancing factor of $\alpha$.
To preserve modality-specific and cross-modality embeddings, MedTok divides the entire $codebook$ into three regions: a text-specific region, a graph-specific region, and a shared region. The shared region includes the graph-shared and text-shared region and is shared with both two modalities. MedTok then queries distinct regions of the $codebook$ to generate their respective tokens and quantized vectors, which are represented by: ($\mathcal{I}(\mathbf{e}_{t}^{s})$, $\mathcal{I}(\mathbf{e}_{g}^{s})$, $\mathcal{I}(\mathbf{e}_{t}^{c})$, $\mathcal{I}(\mathbf{e}_{g}^{c})$) and ($\mathbf{\hat{e}}_{t}^{s}$, $\mathbf{\hat{e}}_{g}^{s}$, $\mathbf{\hat{e}}_{t}^{c}$, $\mathbf{\hat{e}}_{g}^{c}$). The final $codebook$ objective is as follows:
$\mathcal{L}_{vq}=\mathcal{L}(\mathbf{e}_{t}^{s},\mathbf{\hat{e}}_{t}^{s})+\mathcal{L}(\mathbf{e}_{g}^{s},\mathbf{\hat{e}}_{g}^{s})+\mathcal{L}(\mathbf{e}_{t}^{c},\mathbf{\hat{e}}_{t}^{c})+\mathcal{L}(\mathbf{e}_{g}^{c},\mathbf{\hat{e}}_{g}^{c})$ (5)